OpenStudy (anonymous):

limit as x -> pi/2 of tanx

OpenStudy (anonymous):

\[\lim_{x \rightarrow \ {\pi }/ 2 }\tan(x)={\infty}\] 

OpenStudy (anonymous):

what happens when x -> pi/2

OpenStudy (anonymous):

\[[\frac {sinx}{cosx} - \frac 1 {\cos(x)}] \] \[ = \frac {sinx -1}{cosx.} \ Multiply \ \top \ and \ bottom \ by \ sinx \ + \ 1:\]= (sin^2 x - / [cosx ( sinx + 1)] = - \[\frac {\cos^2 x} { [cosx ( sinx + 1)]} \frac { cosx}{ [ sinx + 1]} \ = \] which approaches\[\frac {0}{2} = 0 \]

OpenStudy (anonymous):

1/0 using lopitals rule -cosx/sinx 0/1 answer is 0

OpenStudy (anonymous):

as x tends to0 the equations tends to infinity

OpenStudy (anonymous):

 wow its actually undefined cos from the left its \[x \rightarrow \pi/2^- \rightarrow \infty\] \[x \rightarrow \pi/2^+\rightarrow -\infty\]